OpenStudy (anonymous):

Write as a decimal: 6%. 0.6 0.06 6.0 What is 16% of 40? 3.2 16 6.4 8

OpenStudy (kropot72):

6% = 6/100 = 0.06 40 * (16/100) = 6.4